In an amusement park, a car rolls on a track as shown below. Find the speed (in m/s) of the car at A, B, and C. Note that the work done by the rolling friction is zero since the point at which the rolling friction acts on the tires is momentarily at rest and therefore has a zero displacement.

v=0 84 m 74 m 37 mVa = ? m/s

Vb = ? m/s

Vc = ? m/s

Thanks!
Answer #1

6 By Energy conservation; hy Chf-hi) = ₂ M ( VA-02) 9.84(84-31) = x VA = 30.35 mis mg (84-14) = £ CVB) - 6 VB = 14 mis 0 mg